# mO Revanth[#1]

### v1

@@ C.2 Text Distribution Between AI Models Within the Same Family
C.3 Embedding Visualization and Semantic Cohesion
To delve into deeper semantic alignment, we visualized embeddings generated by unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
@@ C.3 Embedding Visualization and Semantic Cohesion
C.4 Conclusion
The consistency observed across N-grams distributions, text length patterns, and semantic embeddings among Gemini models substantiates our decision to treat each AI family as an author. These models demonstrate coherent writing styles, shared lexical preferences, and tightly clustered semantic representations - hallmarks of unified authorship. Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each AI model family’s writing behavior.
(a) Word length distribution across five AI models.
(b) Character length distribution across five AI models.
@@ C.4 Conclusion
Figure 5: Top 20 most common 3-gram from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
Appendix D Model selection for detector
@@
We first evaluated three popular monolingual models: RoBERTa-base, Flan-T5-base, and e5-base-v2 using the same training and evaluation splits. RoBERTa-base Liu et al. (2019b) and e5-base-v2 Wang et al. (2022) achieved the most balanced trade-off between classification accuracy and regression error (MSE, MAE), while Flan-T5-base Chung et al. (2022) lagged slightly in F1-macro. These results indicated that a stronger encoder backbone yields more robust performance and motivated exploration of multilingual variants for further gains.
@@
Next, we experimented with XLM-RoBERTa-base Conneau et al. (2019) and Multilingual-e5-base Wang et al. (2024a). Both models leverage cross-lingual pretraining, which in our scenario appears to improve representation of diverse linguistic patterns present in FAIDSet. In particular, XLM-RoBERTa-base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
@@
Finally, we integrated contrastive learning via SimCSE Gao et al. (2021) to further refine the sentence embeddings. We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa-base. The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks. Encouraged by this, we adopted the unsupervised SimCSE approach on XLM-RoBERTa-base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
Based on these experiments, we selected the unsupervised SimCSE XLM-RoBERTa-base model for our final system.
@@
Appendix E Ablation study
E.1 The need to use a vector database
@@
E.2 Clustering Algorithm Selection
To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database. Each algorithm was tasked with grouping text samples into human-written, AI-generated, and human-AI collaborative categories, using both known-generator data (held-out from training) and entirely unseen generator data. Performance was measured in terms of accuracy and F1-macro score.
We encoded each text sample using our pretrained sequence-classification model’s penultimate layer, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes. From the results in Table 8, we draw some key findings:
  • Traditional algorithms show reasonable performance on held-out known generators but degrade notably on unseen generators.
  • Fuzzy C-Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k-Means, with smaller degradation on unseen data.
@@ E.2 Clustering Algorithm Selection
Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our full architecture.

### v2

@@ D.2 Text Distribution Between LLMs Within the Same Family
D.3 Embedding Visualization and Semantic Cohesion
To delve into deeper semantic alignment, we visualized embeddings generated by an unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
@@ D.3 Embedding Visualization and Semantic Cohesion
D.4 Conclusion
The consistency observed across N-gram distributions, text length patterns, and semantic embeddings among Gemini models substantiates our decision to treat each LLM family as an author. These models demonstrate coherent writing styles, shared lexical preferences, and tightly clustered semantic representations, hallmarks of unified authorship. Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each LLM family’s writing behavior.
(a) Word length distribution across five LLMs.
(b) Character length distribution across five LLMs.
@@ D.4 Conclusion
Figure 5: Top 20 most common trigrams from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
Appendix E Model Selection for Detector
@@
We first evaluated three popular monolingual models: RoBERTa‑base, Flan‑T5‑base, and e5‑base‑v2 using the same training and evaluation splits. RoBERTa‑base Liu et al. (2019) and e5‑base‑v2 Wang et al. (2022) achieved the most balanced trade‑off between classification accuracy and regression error (MSE, MAE), while Flan‑T5‑base Chung et al. (2022) lagged slightly in F1‑macro. These results indicate that a stronger encoder backbone yields more robust performance, motivating the exploration of multilingual variants for further gains.
@@
Next, we experimented with XLM‑RoBERTa‑base Conneau et al. (2019) and Multilingual‑e5‑base Wang et al. (2024a). Both models leverage cross-lingual pretraining, which, in our scenario, appears to improve the representation of diverse linguistic patterns present in FAIDSet. In particular, XLM‑RoBERTa‑base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
@@
Finally, we integrated contrastive learning via SimCSE Gao et al. (2021) to further refine the sentence embeddings. We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa‑base. The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks. Encouraged by this, we adopted the unsupervised SimCSE approach on XLM‑RoBERTa‑base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
Based on these experiments, we selected the unsupervised SimCSE XLM‑RoBERTa‑base model for our final system.
@@
Appendix F Ablation Study
F.1 The Need to Use a Vector Database
@@
F.2 Clustering Algorithm Selection
To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database. Each algorithm was tasked with grouping text samples into human‐written, AI‐generated, and human–LLM collaborative categories, using both known‐generator data (held out from training) and entirely unseen generator data. Performance was measured in terms of accuracy and F1-macro score.
We encoded each text sample using the penultimate layer of our pre-trained sequence-classification model, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes. From the results in Table 11, we draw some key findings:
  • Traditional algorithms show reasonable performance on held‐out known generators but degrade notably on unseen generators.
  • Fuzzy C‐Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k‐Means, with smaller degradation on unseen data.
@@ F.2 Clustering Algorithm Selection
Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our whole architecture.